An institute is planning on doing a survey to get a 96% confidence interval for a population proportion with a margin of error that is less than 3.5%. They should plan on getting a sample size of at least ______________________.
Answer #1

Solution :

Given that,

\hat p = 1 - \hat p = 0.5

margin of error = E = 0.035

Z\alpha/2 = 2.054

sample size = n = (Z\alpha / 2 / E )2 * \hat p * (1 - \hat p )

= (2.054 / 0.035)2 * 0.5 * 0.5

= 861

sample size = 861